# Ultra-Faint Milky Way Satellites Discovered in Carina, Phoenix, and Telescopium with DELVE Data Release 3

### v1

@@ 6.3 DELVE 7
DELVE 7 is broadly representative of the population of ultra-faint compact satellites discovered in recent years (e.g., Muñoz et al., 2012; Kim et al., 2015; Kim & Jerjen, 2015c; Mau et al., 2019; Cerny et al., 2023a, b; Smith et al., 2024). While past convention has often been to refer to these systems as ultra-faint star clusters, recent investigations into the population of Milky Way satellites at $r_{1/2}<15$ pc have tentatively suggested that some of these systems may in fact be galaxy candidates (Simon et al. 2024; Cerny et al. in prep.). Thus, in the absence of any spectroscopic information about the system’s internal kinematics and metallicity distribution, either a dwarf galaxy or a star cluster classification remains possible for DELVE 7. However, because no galaxies have yet been confirmed in this size or luminosity regime, we conservatively assume that the system is more likely to be a star cluster and name it DELVE 7.
@@ 7 Summary
Carina IV, Phoenix III, and DELVE 7 bring the total number of ultra-faint Milky Way satellites discovered in the DELVE data to 17 (Mau et al. 2020; Cerny et al. 2021a, b, 2023a, 2023b, 2023c, 2024; Tan et al. 2025c). These three candidates were identified using the DELVE DR3 catalog, which provides deeper and more accurate photometric measurements than previous DELVE releases (Drlica-Wagner et al., 2021, 2022). Two of the candidates lie within the DES portion of the catalog, which was constructed from the full six years of DES Y6 observations and therefore provides generally deeper data than the rest of DELVE DR3.
@@
Even with a complete search using data from the current surveys, we expect to be largely incomplete for faint and distant systems such as those presented in this analysis. However, the upcoming Rubin LSST (Ivezić et al., 2019) is expected to discover hundreds of ultra-faint dwarf galaxies in the Local Volume (Hargis et al., 2014; Mutlu-Pakdil et al., 2021; Manwadkar & Kravtsov, 2022; Tsiane et al., 2025). Space-based facilities, including the Roman Space Telescope (Spergel et al., 2015) and Euclid (Euclid Collaboration et al., 2022), are also anticipated to uncover additional ultra-faint dwarfs (Nadler et al., 2024). This rapidly growing census of ultra-faint dwarfs will provide critical insight into the processes of galaxy formation at the smallest scales and serve as a powerful probe of dark matter physics.
@@ Acknowledgment
We would like to thank the Zooniverse team, beta testers, and volunteers for their support in setting up the Zooniverse beta test, and especially citizen scientist and beta tester Ernest Jude P. Tiu, who was the first to identify Phoenix III during the beta test. For this work, we did not make use of the classifications made by other Zooniverse citizen scientists or beta testers. However, we plan to provide a more detailed description of those efforts in a forthcoming paper.
@@
CYT performed the dwarf galaxy search, carried out the ugali morphology fits, produced most of the plots and tables, and led the writing of the paper. WC conducted the analysis of DELVE 7 based on the follow-up IMACS imaging, generated Figure 3, assisted with candidate identification, and contributed to the writing. ABP conducted the CaHK and proper motion analysis, contributed Figure 5, assisted with candidate identification, and contributed with writing. JAS and KO helped identify dwarf galaxy candidates through preparation for the Zooniverse project. ADW provided direct supervision of the research and contributed with writing. JDS carried out the IMACS follow-up observations of DELVE 7, while DJS and BMP reduced the imaging data and produced the catalogs. AMS performed the orbital modeling of Carina IV. DE, PSF, FS internally reviewed the paper. The authors from KRA to AKV contributed to producing and characterizing one or more of the following data products used in the paper: DES Y6 source catalog, DELVE DR3 source catalog, MAGIC catalog, and/or provided valuable comments that improved the paper’s clarity and quality. Builders: The remaining authors contributed to this work through the construction of DECam and other aspects of data collection; data processing and calibration; developing widely used methods, codes, and simulations; running pipelines and validation tests; and promoting the science analysis.
@@ Appendix A Members stars of Carina IV
In Section 5, we discuss the Gaia proper motions and CaHK photometric metallicity measurements of Carina IV candidate member stars, which further confirm its identity as a metal-poor Milky Way satellite. To facilitate spectroscopic follow-up of Carina IV member stars, we include in Table 2 the measured proper motions and metallicities for all stars with $g_{0}<21.3$ and high ugali membership probabilities ($p>0.2$).
Table 2: Properties of stars with high ugali membership probabilities for Carina IV ($p>0.2$) and $g_{0}<21.3$. The stars are divided into three categories: (1) likely members, characterized by consistent Gaia proper motion measurements and low CaHK metallicities (${\rm[Fe/H]_{\rm CaHK}}<-1.5$); (2) possible members, with consistent Gaia proper motions but higher metallicities; and (3) non-members, exhibiting inconsistent proper motions.

### v2

@@ VI.3 DELVE 7
DELVE 7 is broadly representative of the population of ultra-faint compact satellites discovered in recent years (e.g., Muñoz et al., 2012; Kim et al., 2015; Kim and Jerjen, 2015a; Mau et al., 2019; Cerny et al., 2023b, a; Smith et al., 2024). While past convention has often been to refer to these systems as ultra-faint star clusters, recent investigations into the population of Milky Way satellites at $r_{1/2}<15$ pc have tentatively suggested that some of these systems may in fact be galaxy candidates (Simon et al., 2024; Cerny et al., 2026). Thus, in the absence of any spectroscopic information about the system’s internal kinematics and metallicity distribution, either a dwarf galaxy or a star cluster classification remains possible for DELVE 7. However, because no galaxies have yet been confirmed in this size or luminosity regime, we conservatively assume that the system is more likely to be a star cluster and name it DELVE 7.
@@ VII Summary
Carina IV, Phoenix III, and DELVE 7 bring the total number of ultra-faint Milky Way satellites discovered in the DELVE data to 17 (Mau et al. 2020; Cerny et al. 2021a, b, 2023b, 2023a, 2023c, 2024; Tan et al. 2025a). These three candidates were identified using the DELVE DR3 catalog, which provides deeper and more accurate photometric measurements than previous DELVE releases (Drlica-Wagner et al., 2021, 2022). Two of the candidates lie within the DES portion of the catalog, which was constructed from the full six years of DES Y6 observations and therefore provides generally deeper data than the rest of DELVE DR3.
@@
Even with a complete search using data from the current surveys, we expect to be largely incomplete for faint and distant systems such as those presented in this analysis. However, the upcoming Rubin LSST (Ivezić et al., 2019) is expected to discover hundreds of ultra-faint dwarf galaxies in the Local Volume (Hargis et al., 2014; Mutlu-Pakdil et al., 2021; Manwadkar and Kravtsov, 2022; Tsiane et al., 2025). Space-based facilities, including the Roman Space Telescope (Spergel et al., 2015) and Euclid (Euclid Collaboration et al., 2022), are also anticipated to uncover additional ultra-faint dwarfs (Nadler et al., 2024). This rapidly growing census of ultra-faint dwarfs will provide critical insight into the processes of galaxy formation at the smallest scales and serve as a powerful probe of dark matter physics.
@@ Acknowledgment
We thank the anonymous referee for the many useful comments that helped us improve this manuscript. We would also like to thank the Zooniverse team, beta testers, and volunteers for their support in setting up the Zooniverse beta test, and especially citizen scientist and beta tester Ernest Jude P. Tiu, who was the first to identify Phoenix III during the beta test. For this work, we did not make use of the classifications made by other Zooniverse citizen scientists or beta testers. However, we plan to provide a more detailed description of those efforts in a forthcoming paper.
@@
CYT performed the dwarf galaxy search, carried out the ugali morphology fits, produced most of the plots and tables, and led the writing of the paper. WC conducted the analysis of DELVE 7 based on the follow-up IMACS imaging, generated Figure 3, assisted with candidate identification, and contributed to the writing. ABP conducted the CaHK and proper motion analysis, contributed Figure 5, assisted with candidate identification, and contributed with writing. JAS and KO helped identify dwarf galaxy candidates through preparation for the Zooniverse project, and contributed to Figure 1. ADW provided direct supervision of the research and contributed with writing. JDS carried out the IMACS follow-up observations of DELVE 7, while DJS and BMP reduced the imaging data and produced the catalogs. AMS performed the orbital modeling of Carina IV. DE, PSF, FS internally reviewed the paper. The authors from KRA to AKV contributed to producing and characterizing one or more of the following data products used in the paper: DES Y6 source catalog, DELVE DR3 source catalog, MAGIC catalog, and/or provided valuable comments that improved the paper’s clarity and quality. Builders: The remaining authors contributed to this work through the construction of DECam and other aspects of data collection; data processing and calibration; developing widely used methods, codes, and simulations; running pipelines and validation tests; and promoting the science analysis.
@@ Appendix A Members stars of Carina IV
In Section V, we discuss the Gaia proper motions and CaHK photometric metallicity measurements of Carina IV candidate member stars, which further confirm its identity as a metal-poor Milky Way satellite. To facilitate spectroscopic follow-up of Carina IV member stars, we include in Table 2 the measured proper motions and metallicities for all stars with $g_{0}<21.3$ and high ugali membership probabilities ($p_{\rm ugali}>0.2$). For the other two systems, Phoenix III and DELVE 7, as shown in Fig. 2, we find very few stars with high ugali membership probabilities above the Gaia detection limit of $G\sim 21$. For Phoenix III, the only source is the BHB candidate Gaia DR3 4959457079527646080, with $p_{\rm ugali}=0.99$, $g_{0}=20.8$, and $r_{0}=21.0$. This star however does not have a Gaia proper-motion measurement. In the case of DELVE 7, none of the high-probability member stars are detected by Gaia.
Table 2: Properties of stars with high ugali membership probabilities for Carina IV ($p_{\rm ugali}>0.2$) and $g_{0}<21.3$. The stars are divided into three categories: (1) likely members, characterized by consistent Gaia proper motion measurements and low CaHK metallicities (${\rm[Fe/H]_{\rm CaHK}}<-1.5$); (2) possible members, with consistent Gaia proper motions but higher metallicities; and (3) non-members, exhibiting inconsistent proper motions.
@@
Appendix B Phoenix III as a possible intermediate age Star Cluster
As discussed in Section VI.2, the morphological fits for Phoenix III yield two viable solutions. One corresponds to an old, metal-poor stellar population with an age of approximately 12 Gyr, which we argue is the more likely interpretation based on its inclusion of the candidate BHB star Gaia DR3 4959457079527646080. Upcoming HST follow-up observations are expected to confirm or reject this assumption (McQuinn et al., 2025). However, for completeness, we present in the appendix a brief description of the intermediate-age ($\sim$7 Gyr) alternative solution identified by ugali. For this alternative solution, the posterior peaks at a distance modulus of of $(m-M)_{0}=20.5_{-0.1}^{+0.2}$, age of $\tau=7.0_{-1.5}^{+1.7}$ and a metallicity of $Z_{\rm phot}=0.00040_{-0.00016}^{+0.00022}$. In Figure 7, we compare this intermediate-age solution with the old 12 Gyr solution adopted in the main analysis using color–magnitude diagrams in the grg-r band (as in the main analysis) and the additional gig-i band information.
Figure 7: Color–magnitude diagrams for both grg-r (left) and gig-i (right) band pairs for stars within $1\farcm 5$ of Phoenix III. Stars are colored by separation from each system’s centroid, with darker colors representing smaller separations. We overlay PARSEC isochrones corresponding to the intermediate-age 7 Gyr solution (solid maroon lines) and the old 12 Gyr solution (black dotted lines) for comparison. The median photometric uncertainties for stars around Phoenix III as a function of $g$-band magnitude, are indicated by red error bars.